OpenStudy (liliakarina):

About geometric place. To get an equation that satisfies the geometric place of the P(x,y) point, what's the first step? The equation is in the form y=mx+b, but it has to be at equal distance from point (2,1) to (-4,3) ?

OpenStudy (anonymous):

One solution would be any straight line parallel to a line through the two points.

OpenStudy (liliakarina):

@wsgalinaitis I just have to calculate the distance between the points and then the eq?

OpenStudy (anonymous):

So, when I read this question I think I understand what you are asking, however, I will restate it. To begin, Let's solve for the mid point of the segment that has endpoints (2,1) and (-4,3). First, \[Distance = \sqrt{(x_2-x_1)^2-(y_2-y_1)^2}\] Secret hint, the midpoint of two points is the sum of the components and then divided by 2. This means that the midpoint M is \[((-4+2)/2,1+3)/2)=(-1,2)\] Now check that with the distance equation and you will see that they are equidistant apart.

OpenStudy (anonymous):

If you would like I can prove it generally that the midpoint will always be the sum of the components divided by 2, all we do is replace x1 with \[x_1=(x_2+x_1)/2\] And the same with y1. This will give you the original distance over 2 after you take the 4 out of the square root.

OpenStudy (anonymous):

Now, If you want a line passing through the point (2,1) and (-4,3), we need to find the slope, m. This means we use the equation \[m=Rise/Run=(y_2-y_1)/(x_2-x_1)\] and so our slope is -6/2=-3

OpenStudy (anonymous):

just plug this back into the equation y=mx+b and using m=3, (x=2,y=1) and solve for b. and you have your answer.